OpenStudy (anonymous):

What is the distance between points P(10, 7) and Q(1, 8)?

OpenStudy (anonymous):

In the form of a radical.

OpenStudy (er.mohd.amir):

use formula D*D=(x2-x1)^2+(y2-y1)^2 and (x1,y1)=(10,7) so on take root since D*D in LHS

OpenStudy (anonymous):

Isn't that the formula for midpoint?

OpenStudy (whpalmer4):

distance between two points \((x_1,y_1)\) and \((x_2,y_2)\) is given by \[d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] which is really the Pythagorean theorem in disguise :-)

OpenStudy (whpalmer4):